how do i solve this problem x^2 + 10x + 16 = 0

Answer by palanisamy(496)

The given equation is
x^2 + 10x + 16 = 0
x^2 + 2x+ 8x + 16 = 0
x(x+2)+8(x+2) = 0
(x+2)(x+8) = 0
x+2 = 0 or x+8 = 0
x = -2 , -8